Cost figures for a hypothetical firm are given in the following table. Use them for the exercises





below. The firm is selling in a perfectly competitive market.





a. Fill in the blank columns.





b. What is the minimum price needed by the firm to break even?





c. What is the shutdown price?





d. At a price of $40, what output level would the firm produce? What would its profits be?

1
Expert's answer
2022-01-13T08:59:01-0500


b)



"Q=\frac {FC}{p-AVC}""3=\frac {50}{p-27}""p=43.7"

c)



"5=\frac {50}{p-34}""p=44"

d) if p=40, then Q=4



"Profit=TR-TC""TR=pQ=40\times4=160""Profit=160-170=-10"